Phyllis invested $13,500, a portion earning a simple interest rate of 5 1/5% per year and the rest earning a rate of 5% per year. After 1 year the total interest earned on these investments was $692. How much money did she invest at each rate?
Answer by ikleyn(52780)

.
Phyllis invested $13,500, a portion earning a simple interest rate of 5 1/5% per year
and the rest earning a rate of 5% per year.
After 1 year the total interest earned on these investments was $692.
How much money did she invest at each rate?
~~~~~~~~~~~~

Let x be the amount invested at 5 1/5%  (in what follows, I will write 5.2%, for simplicity).

Then the amount invested at 5% is the rest (13500-x) dollars.


As you read the problem, write the total annual interest equation

    0.052x + 0.05*(13500-x) = 692  dollars.


Simplify, express x and get the answer

    0.052x + 0.05*13500 - 0.05x = 692

    0.052x - 0.05x = 692 - 0.05*13500

         0.002x    =      17

              x    =      17/0.002 = 17000/2 = 8500.


ANSWER.  $8500 invested at 5.2%;  the rest  $13500 - $8500 = $5000 invested at 5%.


CHECK.   0.052*8500 + 0.05*5000 = 692 dollars, the total annual interest.  ! Correct !

Solved.